OpenStudy (anonymous):

For the graph e^(sinx) how many zeros are there on the closed interval {0,2pi}

OpenStudy (loser66):

0

OpenStudy (loser66):

\(e^{something} >0 \) for whatever the exponent is

OpenStudy (anonymous):

There is at least one zero on this graph

OpenStudy (loser66):

how??

OpenStudy (loser66):

-1<sin <1

OpenStudy (loser66):

even when sin =0 , e^0 =1>0 no zero

OpenStudy (loser66):

when sin =-1 , \(e^{-1}=\dfrac{1}{e}>0\) still . No zero

OpenStudy (anonymous):

The options are 1,2,3,or 4

OpenStudy (loser66):

If it is so, I am sorry. It is above my head.

OpenStudy (anonymous):

Okay thank you anyway!

OpenStudy (loser66):

https://www.desmos.com/calculator/xwuw3tji5m

OpenStudy (freckles):

there has to be some kinda mistake... like as @Loser66 says exp(sin(x)) is never zero exp( ) function makes sure of that i wonder if they meant the derivaitive's zero though they didn't say this

OpenStudy (freckles):

\[f(x)=e^{\sin(x)} \\ f'(x)=\cos(x)e^{\sin(x)} \\ \cos(x)e^{\sin(x)}=0 \\ \\\ \text{ here you would only need to solve } \cos(x)=0 \text{ on } [0,2\pi]\] now again this is the zeros of your first derivative exp(sin(x)) doesn't have any zeros itself